a rectangle has a perimeter of 320 meters and its length is 3 times the width. Find the dimensions of the rectangle
Answer by addingup(3677)

2Length+2Width = Perimeter
2L+2W = 320
and
L = 3W
---------------------------
2(3W)+2W = 320
6W+2W = 320
8W = 320
W = 40
and
L = 3(40) = 120
---------------------------
Check:
2(120)+2(40) = 320
240+80 = 320 Correct
:
John